# Confinement-induced collective motion in suspensions of run-and-tumble particles

### v1

@@ Abstract
Collective motion is ubiquitous in active systems at all length and time scales. The mechanisms behind such collective motion usually are alignment interactions between active particles, effective alignment after collisions between agents or symmetry-breaking fluctuations induced by passive species in active suspensions. In this article, we introduce a new type of collective motion in the shape of a traveling band induced purely by confinement, where no explicit or effective alignment are prescribed among active agents. We study a suspension of run-and-tumble particles confined in microchannels comprising asymmetric boundaries: one flat wall and one array of funnel-like obstacles. We study the phase behavior of the confined active suspension upon changes in the packing fraction and the persistence length to define the stability region of the traveling band. We characterize the traveling band structurally and dynamically and study its stability with respect to the tilt angle of the obstacles. Lastly, we describe the mechanism of motion of the band, which resembles the tracked locomotion of some heavy vehicles like tractors, finding that a counter-flux of active particles in the lower part of the band, explained in terms of source-sink and vacancy diffusion mechanisms, is the facilitator of the traveling band and sustains its motion. We name this new collective phenomenon confinement-induced tracked locomotion.
@@ I Introduction
Collective motion is ubiquitous in Nature at different length and time scales vicsek2012collective, ranging from genetic material migration during cell division scholey2003cell, to bacterial swarming kearns2010field; darnton2010dynamics, the mesmerizing collective migration of Dictyosteliumhashimura2019collective, the phototactic response of the green micro-algae C. reinhardtii witman1993chlamydomonas; arrieta2017phototaxis; choudhary2019reentrant; choudhary2025transient, swarms of insects topaz2008model; ariel2015locust, schools of fish niwa1994self; lopez2012behavioural; liu2025collective, flocks of birds bialek2012statistical; cavagna2014bird, herds of sheep garcimartin2015flow, and crowds of people gu2025emergence, among other examples.
@@
Thirty years ago, Vicsek and collaborators proposed a model to mimic the collective behavior of aligning active particles: a simple aligning mechanism between point-like agents moving in a noisy environment, affecting the agents’ orientation vicsek1995novel. In particular, at intermediate densities and noise intensities, the Vicsek model predicted the formation of traveling bands (dense regions of aligning particles)chate2008collective; chate2008modeling; solon2015pattern; ginelli2016physics. The alignment between particles was the underlying mechanism of such traveling bands. The effects of confinement on the collective motion of active particles with alignment interactions have recently been addressed. Aligning self-propelled particles confined in arrays of obstacles with different shapes and distributions has revealed trapping and sub-diffusive behavior chepizhko2013diffusion; chepizhko2015active. On the other hand, confinement in ordered arrays of obstacles can induce a re-entrant collective behavior, leading from self-organization into traveling bands to a disordered state and, as the packing fraction of obstacles increases, to a new synchronous collective behavior serna2025influence.
@@
Although the Vicsek model can qualitatively explain many collective phenomena observed in organisms with evolved sensory organs and complex signal processing systems, it cannot do the same for their microscopic counterparts. This is due to colloids or motile microorganisms usually exhibiting collective motion as a response to light or chemical stimuli budrene1995dynamics; wu2015collective; erban2004individual; liebchen2018synthetic; keller1971traveling; arrieta2017phototaxis; choudhary2019reentrant; choudhary2025transient or due to hydrodynamic interactions alarcon2013spontaneous; alarcon2017morphology; zottl2014hydrodynamics, rather than through an explicit alignment mechanism among agents. Thus, systems exhibiting collective motion that arises without aligning interactions between agents and without the action of external fields are highly appreciated in the active matter community and provide a platform for understanding the physics behind this emergence.
@@
Some examples of systems with no explicit aligning interactions that exhibit collective motion in bulk are "energy-depot" active disks which form traveling bands as a consequence of an effective alignment mechanism mediated by collisions miranda2025collective, and active-passive mixtures of Brownian particles that form propagating interfaces as a result of a complex combined mechanism involving source/sink effects and oriented displacements of particles modulated by density fluctuations induced by the presence of passive particles fernandez2025dynamics; wysocki2016propagating.
@@
Geometrical confinement has proven to be an efficient mechanism in rectifying anand2024transport; tailleur2009sedimentation; wan2008rectification; di2010bacterial, trapping and sorting galajda2007wall; martinez2020trapping; kaiser2012capture; kumar2019trapping active particles, especially configurations comprising concave regions in which the active particles tend to accumulate, such as funnels or chevrons reichhardt2017ratchet. Moreover, the shape of the confining boundaries may induce jamming statespajger2025jamming or destroy phase separationben2022disordered in active systems. In this article, we present a simulation study of a two-dimensional suspension of run-and-tumble particles confined in asymmetric microchannels with one flat wall and one wall made of funnel-like obstacles (see Fig. 1).
@@
To the best of our knowledge, we report for the first time the emergence of a dense traveling band whose formation and motion are completely induced and sustained by confinement, with neither explicit alignment between agents, nor effective alignment induced by particle shape or collisions, nor by the presence of passive particles in the suspension. We discovered such a traveling structure serendipitously, while performing the analysis of the results published in Ref.serna2025sorting by some of the present authors. The mechanism of the band’s motion resembles the tracked locomotion of some heavy vehicles such as tractors, and thus we have named it confinement-induced tracked locomotion.
@@ II.1 Model
The system under consideration is a two-dimensional suspension of active particles (type 1, in pink in Fig. 1) following run-and-tumble dynamics confined in microchannels composed of frozen particles (type 2, in yellow in Fig. 1).
Figure 1: Scheme of the geometry of the system. The top boundary of the microchannel is a flat wall, and the bottom boundary of an array of funnel-like obstacles with period $l_{obs}=2\left(L_{obs}\cos\theta_{obs}+d+\sigma_{22}\right)$, where $d=1.25\sigma_{22}$ is the gap size, $L_{obs}=3.0\sigma_{22}$ is the length of the obstacles and $\theta_{obs}$ the angle formed between the obstacles and the $x$-axis. The walls of the microchannel are composed of frozen particles (in yellow) of diameter $\sigma_{22}$, with a lattice constant, $l_{w}=0.12\sigma_{22}$. The axis $y=0$ is placed at the center of the obstacles. The active particles (i pink) follow the run-and-tumble phenotype observed in different motile microorganisms.
@@
Here, $\bm{r}_{i}$, $\bm{n}_{i}$ and $\bm{v}_{i}$ are the position vector, unit orientation vector and velocity vector of particle $i$, respectively. The active self-propulsion force is set to $F_{a}=10$, the friction coefficient, $\gamma=10$, the mass of the particles $m_{i}=1.0$, and the thermal energy, $k_{B}T=1.0$. The self-propulsion speed of the particles is, thus $v_{0}=F_{a}/\gamma=1.0$. With such a choice of parameters, the inertia of the particles’ motion is only significant at very small length scales, so we can consider the system follows an overdamped regimeserna2025sorting. The thermal bath applies a stochastic force, $\bm{W}_{i}=(W^{x}_{i},W^{y}_{i})$, on the $i$th particle, given by the standard Gaussian white noise with zero mean and correlation $\langle W_{i}^{x}(t)W^{y}_{i}(t)\rangle=\delta_{ij}\delta_{xy}\delta(t-t^{\prime})$.
All particles interact via a WCAweeks1971role repulsive potential,
$U(r)=\begin{cases}4\epsilon_{mn}\left[\left(\frac{\sigma_{mn}}{r}\right)^{12}-\left(\frac{\sigma_{mn}}{r}\right)^{6}\right]+\epsilon_{mn},&r<2^{1/6}\sigma_{mn},\\ 0&r\geq 2^{1/6}\sigma_{mn}.\end{cases}$ (2)
with indexes $m$ and $n$ running over the two particle types. We set $\sigma_{11}=2$, $\sigma_{22}=1$ and $\epsilon_{11}=\epsilon_{22}=1$. The crossed interactions are given by the Lorentz-Berthelot mixing ruleslorentz1881ueber; berthelot1898melange: $\sigma_{mn}=(\sigma_{mm}+\sigma_{nn})/2$ and $\epsilon_{mn}=\sqrt{\epsilon_{mm}\epsilon_{nn}}$. Particles constituting the microchannel walls are placed at a distance $l_{w}=0.12\sigma_{22}$ from each other.
@@
$\theta_{i}(t+dt)=\begin{cases}\theta_{i}(t),&\alpha<\zeta{(t)}\\ \theta_{i}(t)+\Phi{(t)}&\alpha>\zeta{(t)}\end{cases},$ (3)
@@
being $\theta_{i}$ the angle formed between the orientation vector $\textbf{n}_{i}$ and the positive $x$-axis. $\alpha=t_{trial}/\tau_{r}$ is the probability of occurrence of a tumbling event, $t_{trial}$ being the time between trials, and $\tau_{r}$ the mean time between tumbling events or reorientation time. Here we use $t_{trial}=dt$, i.e. tumbling trials are performed at every step. Thus, the tumbling rate is $\alpha_{r}=\alpha/t_{trial}=1/\tau_{r}$. $\zeta\in[0,1]$ is a uniformly distributed random number generated at each step and $\Phi\in[0,2\pi]$ is a uniformly distributed random angle. Note that, in this formulation the tumbling rate, $\alpha$, is the probability that a particle performs a tumbling in the current timestep, and thus it is dimensionless.
@@
The particles’ level of activity is thus characterized by the persistence length, $l_{p}=\frac{F_{a}}{\alpha_{r}\gamma}$, which is the average distance an active particle travels before tumbling. In our work, we vary the suspension’s packing fraction, $\phi=\frac{N\pi\sigma_{11}^{2}}{4L_{x}L_{y}}\in[0.20,0.40]$, with $N$ the number of active particles and $L_{x}$ and $L_{y}$ the edge lengths of the rectangular simulation box. We fix $L_{y}=20\sigma_{22}$. The system contains $N\in(1000,20000)$ active particles, depending on the packing fraction and the system’s size, thus $L_{x}$ is chosen accordingly. The initial condition is defined as a random distribution of active particles inside the microchannel with a uniform random distribution of orientations. Periodic boundary conditions are imposed on the $x$-axis. The confining channel comprises funnel-like obstacles in the lower part and a flat wall in the upper part (See Fig. 1). The obstacles feature gaps that do not allow the passage of the active particles. The gaps provide two stable resting positions for the active particles, since they can fit in the gaps (but cannot go through them). In Fig. 1, we present the geometry of the funnel-like obstacles: we set their length to $L_{obs}=3.0\sigma_{22}$, forming an angle $\theta_{obs}$ with respect to the $x$-axis which will be varied in the range $7.5^{o},52.5^{o}]$ in steps of $7.5^{o}$. The width of the gaps is fixed to $d=1.25\sigma_{22}$. The period of the obstacle array is given by $l_{obs}=2\left(L_{obs}\cos\theta_{obs}+d+\sigma_{22}\right)$. As a comparison, we also consider two different channels: one with two flat parallel walls and another consisting of a flat wall and corrugations with no gaps (See Fig S1 of Supplementary Information).
@@
All quantities in the manuscript are expressed in reduced simulation units, using $k_{B}T$, $\sigma_{22}$, and $m$ as units of energy, distance and mass, respectively. The simulations are performed using the open-source Molecular Dynamics package LAMMPSplimpton1995fast; LAMMPS, in which we have implemented the run-and-tumble dynamics in equation 3. The equations of motion are integrated using the NVE integrator and the Langevin’s thermostat with $dt=0.001\tau$, where $\tau=\sqrt{\frac{m\sigma_{22}^{2}}{k_{B}T}}$ is the simulation time unit. Simulations are run for up to $30000\tau$.
@@ II.2.1 Identifying dense structures: Percolation
To start with, we characterize the internal structure of the system and, by analyzing the instantaneous particle configurations, we identify structures such as travelling bands and clogs: the former consist of a large percolating clusters moving along the channel, while the latter are large percolating immobile clusters. For each configuration, the set of pairwise particle distances was evaluated under periodic boundary conditions along the longitudinal direction of the channel (x-axis), and fixed boundaries along the y-axis. Two particles were considered connected if their separation was smaller than a prescribed cutoff distance $r_{\mathrm{cut}}=\sigma_{11}$, chosen slightly above the mean interparticle spacing within the dense phase as shown in the radial distribution function (see Fig. S2, Supplementary Information). From this connectivity criterion, a percolation network was constructed, where each particle represents a node and pairwise connections define edges. Clusters were then identified as the connected components of this network and the system is considered in a percolated state when the network connects particles from one extreme $y=0$ to the other $y=L_{y}-\sigma_{12}$ along y-axis. Having established the largest connected cluster in the system, this will correspond to the traveling band or the clog case for the following analysis. This procedure provides a robust identification of the collective structure and allows one to extract statistical quantities such as the temporal evolution of the band size and the distribution of orientation angles within the dense region.
II.2.2 Following the structures: Band velocity

### v2

@@ Abstract
Collective motion is ubiquitous in active systems at all length and time scales. The mechanisms behind such collective motion usually are alignment interactions between active particles, effective alignment after collisions between agents or symmetry-breaking fluctuations induced by passive species in active suspensions. In this article, we introduce a new type of collective motion in the shape of a traveling band induced purely by confinement, where no explicit or effective alignment are prescribed among active agents. We study a suspension of run-and-tumble particles confined in microchannels comprising asymmetric boundaries: one flat wall and one array of funnel-like obstacles. We study the phase behavior of the confined active suspension upon changes in the packing fraction and the persistence length to define the stability region of the traveling band. We characterize the traveling band structurally and dynamically and study its stability with respect to the geometry of the microchannel. Lastly, we describe the mechanism of motion of the band, which resembles the tracked locomotion of some heavy vehicles like tractors, finding that a counter-flux of active particles in the lower part of the band, explained in terms of source-sink and vacancy diffusion mechanisms, is the facilitator of the traveling band and sustains its motion. We name this new collective phenomenon confinement-induced tracked locomotion.
@@ I Introduction
Collective motion is ubiquitous in Nature at different length and time scales Vicsek and Zafeiris (2012), ranging from genetic material migration during cell division Scholey et al. (2003), to bacterial swarming Kearns (2010); Darnton et al. (2010), the mesmerizing collective migration of DictyosteliumHashimura et al. (2019), the phototactic response of the green micro-algae C. reinhardtii Witman (1993); Arrieta et al. (2017); Choudhary et al. (2019, 2025), swarms of insects Topaz et al. (2008); Ariel and Ayali (2015), schools of fish Niwa (1994); Lopez et al. (2012); Liu et al. (2025), flocks of birds Bialek et al. (2012); Cavagna and Giardina (2014), herds of sheep Garcimartín et al. (2015), and crowds of people Gu et al. (2025), among other examples.
@@
Thirty years ago, Vicsek and collaborators proposed a model to mimic the collective behavior of aligning active particles: a simple aligning mechanism between point-like agents moving in a noisy environment, affecting the agents’ orientation Vicsek et al. (1995). In particular, at intermediate densities and noise intensities, the Vicsek model predicted the formation of traveling bands (dense regions of aligning particles)Chaté et al. (2008b, a); Solon et al. (2015); Ginelli (2016). The alignment between particles was the underlying mechanism of such traveling bands. The effects of confinement on the collective motion of active particles with alignment interactions have recently been addressed. Aligning self-propelled particles confined in arrays of obstacles with different shapes and distributions has revealed trapping and sub-diffusive behavior Chepizhko and Peruani (2013, 2015). On the other hand, confinement in ordered arrays of obstacles can induce a re-entrant collective behavior, leading from self-organization into traveling bands to a disordered state and, as the packing fraction of obstacles increases, to a new synchronous collective behavior Serna et al. (2025b).
@@
Although the Vicsek model can qualitatively explain many collective phenomena observed in organisms with evolved sensory organs and complex signal processing systems, it cannot do the same for their microscopic counterparts. This is due to colloids or motile microorganisms usually exhibiting collective motion as a response to light or chemical stimuli Budrene and Berg (1995); Wu (2015); Erban and Othmer (2004); Liebchen and Lowen (2018); Keller and Segel (1971); Arrieta et al. (2017); Choudhary et al. (2019, 2025) or due to hydrodynamic interactions Alarcón and Pagonabarraga (2013); Alarcón et al. (2017); Zöttl and Stark (2014), rather than through an explicit alignment mechanism among agents. Thus, systems exhibiting collective motion that arises without aligning interactions between agents and without the action of external fields are highly appreciated in the active matter community and provide a platform for understanding the physics behind this emergence.
@@
Some examples of systems with no explicit aligning interactions that exhibit collective motion in bulk are "energy-depot" active disks which form traveling bands as a consequence of an effective alignment mechanism mediated by collisions Miranda et al. (2025), and active-passive mixtures of Brownian particles that form propagating interfaces as a result of a complex combined mechanism involving source/sink effects and oriented displacements of particles modulated by density fluctuations induced by the presence of passive particles Fernandez-Quevedo et al. (2025); Wysocki et al. (2016).
@@
Geometrical confinement has proven to be an efficient mechanism in rectifying Anand et al. (2024); Tailleur and Cates (2009); Wan et al. (2008); Di Leonardo et al. (2010), trapping and sorting Galajda et al. (2007); Martinez et al. (2020); Kaiser et al. (2012); Kumar et al. (2019) active particles, especially configurations comprising concave regions in which the active particles tend to accumulate, such as funnels or chevrons Reichhardt and Reichhardt (2017). Moreover, the shape of the confining boundaries may induce jamming statesPajger and Slanina (2025) or destroy phase separationBen Dor et al. (2022) in active systems. In this article, we present a simulation study of a two-dimensional suspension of run-and-tumble particles confined in asymmetric microchannels with one flat wall and one wall made of funnel-like obstacles (see Fig. 1).
@@
To the best of our knowledge, we report for the first time the emergence of a dense traveling band whose formation and motion are completely induced and sustained by confinement, with neither explicit alignment between agents, nor effective alignment induced by particle shape or collisions, nor by the presence of passive particles in the suspension. We discovered such a traveling structure serendipitously, while performing the analysis of the results published in Ref.Serna et al. (2025a) by some of the present authors. The mechanism of the band’s motion resembles the tracked locomotion of some heavy vehicles such as tractors, and thus we have named it confinement-induced tracked locomotion.
@@ II.1 Model
The system under consideration is a two-dimensional suspension of active particles (type 1, in pink in Fig. 1) following run-and-tumble dynamics confined in microchannels composed of frozen particles (type 2, in yellow in Fig. 1). Figure 1 shows a scheme of the system with a graphic explanation of the geometrical parameters of the channel.
Figure 1: Scheme of the geometry of the system. The top boundary of the microchannel is a flat wall, and the bottom boundary of an array of funnel-like obstacles with period $l_{obs}=2\left(L_{obs}\cos\theta_{obs}+d+\sigma_{22}\right)$, where $d=1.25\sigma_{22}$ is the gap size, $L_{obs}=3.0\sigma_{22}$ is the length of the obstacles and $\theta_{obs}$ the angle formed between the obstacles and the $x$-axis. The walls of the microchannel are composed of frozen particles (in yellow) of diameter $\sigma_{22}$, with a lattice constant, $l_{w}=0.12\sigma_{22}$. The axis $y=0$ is placed at the center of the obstacles. The active particles (in pink) follow the run-and-tumble phenotype observed in different motile microorganisms.
@@
Here, $\bm{r}_{i}$, $\bm{n}_{i}$ and $\bm{v}_{i}$ are the position vector, unit orientation vector and velocity vector of particle $i$, respectively. The active self-propulsion force is set to $F_{a}=10$, the friction coefficient, $\gamma=10$, the mass of the particles $m_{i}=1.0$, and the thermal energy, $k_{B}T=1.0$. The self-propulsion speed of the particles is, thus $v_{0}=F_{a}/\gamma=1.0$. With such a choice of parameters, the inertia of the particles’ motion is only significant at very short time scales, so we can consider the system follows overdamped dynamics Serna et al. (2025a). The thermal bath applies a stochastic force, $\bm{W}_{i}=(W^{x}_{i},W^{y}_{i})$, on the $i$th particle, given by the standard Gaussian white noise with zero mean and correlation $\langle W_{i}^{x}(t)W^{y}_{i}(t)\rangle=\delta_{ij}\delta_{xy}\delta(t-t^{\prime})$.
All particles interact via a WCAWeeks et al. (1971) repulsive potential,
$U(r)=\cases{4}\epsilon_{mn}\left[\left(\frac{\sigma_{mn}}{r}\right)^{12}-\left(\frac{\sigma_{mn}}{r}\right)^{6}\right]+\epsilon_{mn},&r<2^{1/6}\sigma_{mn},\\ 0&r\geq 2^{1/6}\sigma_{mn}.$ (2)
with indexes $m$ and $n$ running over the two particle types (1: active particles, 2: wall particles). We set $\sigma_{11}=2$, $\sigma_{22}=1$ and $\epsilon_{11}=\epsilon_{22}=1$. The crossed interactions are given by the Lorentz-Berthelot mixing rulesLorentz (1881); Berthelot (1898): $\sigma_{mn}=(\sigma_{mm}+\sigma_{nn})/2$ and $\epsilon_{mn}=\sqrt{\epsilon_{mm}\epsilon_{nn}}$. Particles constituting the microchannel walls are placed at a distance $l_{w}=0.12\sigma_{22}$ from each other. Note that wall particles (type 2) do not interact with each other, they only interact with active particles (type 1).
@@
$\theta_{i}(t+dt)=\cases{\theta}_{i}(t),&\alpha<\zeta{(t)}\\ \theta_{i}(t)+\Phi{(t)}&\alpha>\zeta{(t)},$ (3)
@@
being $\theta_{i}$ the angle formed between the orientation vector $\textbf{n}_{i}$ and the positive $x$-axis. $\alpha=t_{trial}/\tau_{r}$ is the probability of occurrence of a tumbling event, $t_{trial}$ being the time between trials, and $\tau_{r}$ the mean time between tumbling events or reorientation time. Here we use $t_{trial}=dt$, i.e. tumbling trials are performed at every step. Thus, the tumbling rate is $\alpha_{r}=\alpha/t_{trial}=1/\tau_{r}$. $\zeta\in[0,1]$ is a uniformly distributed random number generated at each step and $\Phi\in[0,2\pi]$ is a uniformly distributed random angle.
@@
The particles’ level of activity is thus characterized by the persistence length, $l_{p}=\frac{F_{a}}{\alpha_{r}\gamma}$, which is the average distance an active particle travels before tumbling. In our work, we vary the suspension’s packing fraction, $\phi=\frac{N\pi\sigma_{11}^{2}}{4L_{x}L_{y}}\in[0.20,0.40]$, with $N$ the number of active particles and $L_{x}$ and $L_{y}$ the edge lengths of the rectangular simulation box. We fix $L_{y}=10\sigma_{11}$ for most of the simulations. We will state it explicitly when chosen otherwise. The system contains $1000\leq N\leq 20000$ active particles, depending on packing fraction and system size. $L_{x}$ is chosen and adjusted accordingly. The initial condition is defined as a random distribution of active particles inside the microchannel with a uniform random distribution of orientations. Periodic boundary conditions are imposed on the $x$-axis. The confining channel comprises funnel-like obstacles in the lower part and a flat wall in the upper part (see Fig. 1). The obstacles feature gaps that do not allow the passage of the active particles. The gaps provide two stable resting positions for the active particles, since they can fit in the gaps (but cannot go through them). In Fig. 1, we present the geometry of the funnel-like obstacles: we set their length to $L_{obs}=3.0\sigma_{22}$, forming an angle $\theta_{obs}$ with respect to the $x$-axis which will be varied in the range $0^{o},75^{o}]$. The width of the gaps is fixed to $d=1.25\sigma_{22}$. The period of the obstacle array is given by $l_{obs}=2\left(L_{obs}\cos\theta_{obs}+d+\sigma_{22}\right)$. For comparison, we also consider three different channel types: the first with two flat parallel walls, the second consisting of a top flat wall and a bottom wall made of corrugations with no gaps (see Fig. S1 of Supplementary Information) , and the third composed of a top wall made of trapezoids and a bottom wall comprising the original funnel-like obstacles with gaps (see Fig.6b) .
@@
All quantities in the manuscript are expressed in reduced simulation units, using $k_{B}T$, $\sigma_{22}$, and $m$ as units of energy, distance and mass, respectively. The simulations are performed using the open-source Molecular Dynamics package LAMMPSPlimpton (1995); Thompson et al. (2022), in which we have implemented the run-and-tumble dynamics in equation 3. The equations of motion are integrated using the NVE integrator and the Langevin’s thermostat with $dt=0.001\tau$, where $\tau=\sqrt{\frac{m\sigma_{22}^{2}}{k_{B}T}}$ is the simulation time unit. Most of the simulations are run for up to $30000\tau$, but some are run for longer to check the stability of some of the observed self-organized steady states.
@@ II.2.1 Identifying dense structures: Percolation
To start with, we characterize the internal structure of the system and, by analyzing the instantaneous particle configurations, we identify structures such as traveling bands and clogs: the former consist of a large percolating cluster moving along the channel, while the latter consists of a large percolating immobile cluster. For each configuration, the set of pairwise particle distances was evaluated under periodic boundary conditions along the longitudinal direction of the channel (x-axis) and fixed boundaries along the y-axis. Two particles were considered connected if their separation was smaller than a prescribed cutoff distance $r_{\mathrm{cut}}=\sigma_{11}$, chosen slightly above the mean interparticle spacing within the dense phase as shown in the radial distribution function (see Fig. S2, Supplementary Information). From this connectivity criterion, a percolation network was constructed, where each particle represents a node and pairwise connections define edges. Clusters were then identified as the connected components of this network, and the system is considered in a percolated state when the network connects particles from one extreme, $y=0$, to the other, $y=L_{y}-\sigma_{12}$, along the y-axis. Having established the largest connected cluster in the system, this will correspond to the traveling band or the clog case for the following analysis. This procedure provides a robust identification of the collective structure. It allows the calculation of different observables, such as the temporal evolution of the band size and the distribution of orientation angles within the dense region.
II.2.2 Following the structures: Band’s velocity